OpenStudy (anonymous):

A pulley is moved by a belt and turns trough an angle of 420degrees while a point on the belt moves 16 feet. Find the radius of the pulley to the nearest foot

OpenStudy (anonymous):

@RadEn This is my last problem, and i have no idea, if its not too much trouble, can i get one last bit of help?

OpenStudy (e.mccormick):

Did you try and sketch it? That can help you know what values go where.

OpenStudy (anonymous):

I guess i'l try that, do you know how to do it though?

OpenStudy (e.mccormick):

Yes.

OpenStudy (e.mccormick):

Know the relationship between revolutions and degrees?

OpenStudy (anonymous):

360 in one revolution

OpenStudy (e.mccormick):

Good. Now, know the circumference formula?

OpenStudy (anonymous):

2 pi R

OpenStudy (anonymous):

how would i connect circumference and degrees here?

OpenStudy (e.mccormick):

Ah, the circumfence is the distance of one revolution, or 360 degrees.

OpenStudy (e.mccormick):

So you know what one revolution is, the formula for the circumference, the angle moved, and the measure moved for that angle. Can you set up a ratio with those four things?

OpenStudy (anonymous):

I'm not seeing how the 16 feet is going to play in here

OpenStudy (e.mccormick):

It is a distance, just like the circumference is a distance. So you have two distance measures and two degree measures. Each distance is related to a degrees: \(\dfrac{\text{Distance}}{\text{Degrees}}\)

OpenStudy (anonymous):

so 16/420, thats half the ratio

OpenStudy (e.mccormick):

Yes. And the other?

OpenStudy (anonymous):

the radius/360?

OpenStudy (e.mccormick):

Not the radius. The circumference.

OpenStudy (anonymous):

ooooh

OpenStudy (e.mccormick):

\(\dfrac{16}{420}=\dfrac{2\pi r}{360}\) Then solve for r. \(\ddot \smile\)

OpenStudy (anonymous):

i get that, but what am i going to do with these two

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

so, 16/420, then times 360, then divide by 2 and pi?

OpenStudy (e.mccormick):

See, you are measuring the distance twice on the same circle. That is why this works. Because they are the same circle, they MUST have an equal ratio.

OpenStudy (e.mccormick):

That would work. I might simplify before I did the rest, but it is all personal choice at that point.

OpenStudy (anonymous):

all right, thanks :D

OpenStudy (e.mccormick):

Where a sketch can help in this situation is it would let you see that you had one pulley and the distance around it. That let me know what I had to work with. Everything else is just math principals after that. Have fun!

OpenStudy (anonymous):

yeah, a whole sheet of this same stuff.. its gonna be a blast!